### v1

# MACI: Multi-Agent Collaborative Intelligence for Robust Reasoning and Temporal Planning

Abstract
Artificial intelligence requires deliberate reasoning, temporal awareness, and effective constraint management—capabilities beyond the pattern-matching strengths of LLMs. LLMs struggle with planning tasks because of their reliance on associative reasoning, inability to self-verify, and inconsistent constraint awareness. We propose Multi-Agent Collaborative Intelligence (MACI), a framework centered on a meta-planner (MP) that orchestrates multiple agents to generate planner templates that define roles and constraints. These planners produce actionable workflows of role nodes and dependency constraints, enabling advanced temporal reasoning and adaptability.
MACI’s three-tier architecture includes a meta-planning module for planner construction, common agents for general reasoning, and specialized agents for domain expertise. By decoupling planning from validation, it overcomes key LLM limitations. Evaluations demonstrate MACI’s effective constraint satisfaction, conflict detection, and reasoning, positioning it as a robust solution for complex reasoning and planning tasks.
Machine Learning, ICML
1 Introduction
The development of artificial general intelligence requires a shift from pattern matching to deliberate analytical reasoning. Large Language Models (LLMs) excel at pattern matching but lack deliberative reasoning required for planning tasks (Kahneman, 2011). These capabilities become crucial for tasks requiring sustained attention, awareness of constraints, and reasoning about future states.
1.1 Limitations of LLMs in Planning
Our evaluation reveals three critical limitations in current LLMs that prevent effective planning.
Metacognitive Limitations in Self-verification
LLM limitations in self-verification extend beyond Gödel’s incompleteness theorems for formal systems (Gödel, 1967). Though formal systems cannot prove their own consistency within their framework, LLMs face greater challenges due to their probabilistic nature and lack of logical foundations (detailed arguments in Section 2). This limitation provides theoretical justification for distributed validation, where specialized agents perform independent verification of planning processes.
Attention-based Processing Bias
In complex planning scenarios, LLMs demonstrate a form of cognitive tunneling where recent context dominates earlier-established constraints. For example, when planning a multi-stop journey, an LLM might focus on optimizing the final leg’s arrival time while losing sight of earlier constraints about vehicle availability or required rest periods. This attention bias leads to locally optimized but globally invalid solutions, undermining the overall feasibility of the plan. The inability to maintain consistent awareness of all constraints throughout the planning process represents a fundamental limitation in current architectures.
Common Sense Knowledge Integration
LLMs often overlook practical constraints that humans naturally consider. In travel planning, they might ignore factors such as airport security wait time or luggage claim delays. These omissions are the result of an inadequate integration of common sense knowledge into the planning process. Although humans automatically incorporate such practical considerations, LLMs require explicit specification of these constraints, often leading to incomplete or impractical plans.
Table 1: Comparison of Validation Limitations and Solutions
Gödel’s Incomp. Theorems LLM Validation Limitations MACI’s Solution
A formal system cannot prove its own consistency LLMs cannot reliably validate their own outputs due to probabilistic foundations MACI uses external validation agents to bypass self-referential loops
Consistency proofs require a “higher” system Validation requires external frameworks (e.g., logic engines, human oversight) Distributed validation acts as the “higher" system for LLM outputs
Self-reference leads to paradox or incompleteness LLMs hallucinate or contradict themselves when critiquing their own outputs Meta-planning separates planning and validation into distinct agent roles
1.2 The MACI Framework
To overcome the inherent limitations of single-model approaches in planning tasks, we introduce the Multi-Agent Collaborative Intelligence (MACI) framework. At the core of MACI lies the meta-planner (MP), a high-level orchestration mechanism that collaborates with multiple agents to construct a specialized planner for temporal reasoning and scheduling. By enabling agents to work collectively at the meta-level, MACI ensures robust and adaptive planning across complex domains.
Meta-Planning Module
The meta-planning module serves as the cornerstone of MACI, analyzing task requirements and dynamically generating a planner specifically tailored to address complex challenges. This planner, in turn, produces a workflow composed of nodes representing roles (e.g., cook, driver, supervisor) and dependencies representing constraints (e.g., temporal, spatial, resource). Unlike traditional multi-agent systems, MACI’s meta-planner orchestrates a dynamic network of agents, validates their interactions, and enforces protocols to ensure comprehensive planning and execution. This layered approach—planning to generate a planner that produces actionable workflows—positions MACI as a forward-looking solution to real-world planning problems.
Common Agents
Common agents form the foundation of MACI’s general-purpose planning capabilities. They address critical tasks such as practical reasoning, constraint validation, and robustness analysis. For example, the Common Sense Integration agent identifies implicit constraints, the Constraint Validation agent ensures solution feasibility, and the Plan Evaluation agent measures performance metrics. These agents collectively maintain both explicit and implicit constraints while ensuring compliance and safety across domains.
Specialized Agents
Specialized agents extend MACI’s capabilities by addressing domain-specific requirements. Instead of overburdening a single planner, MACI deploys focused agents for tasks such as travel scheduling, financial risk assessment, or logistics optimization. These agents integrate domain expertise with the foundational work of common agents, ensuring precise and effective solutions for specialized planning challenges.
1.3 Contributions
This work introduces MACI as a novel approach to collaborative planning. Preliminary evaluations in travel planning scenarios demonstrate MACI’s superiority in constraint satisfaction, conflict detection, and practical reasoning compared to single-LLM approaches. Key contributions of this paper include:
  1. 1.
    Identifying limitations of single-LLM approaches through systematic experimentation.
  2. 2.
    Developing a three-tier architecture for distributed planning and validation.
  3. 3.
    Designing MACI with common and specialized agents for constraint management and reasoning.
  4. 4.
    Demonstrating MACI’s effectiveness in travel planning scenarios, achieving significant improvements over single-LLM systems.
The remainder of this paper is organized into five sections. Section 2 reviews related MAS frameworks and explains why they fall short for the task at hand. Section 3 examines the limitations of LLMs through case studies. Section 4 details the specifications of the MACI framework. Section 5 evaluates MACI’s performance in comparison to state-of-the-art LLMs. Finally, Section 6 discusses implications for future research and alternative perspectives.
2 Related Work
The development of MACI builds on theoretical insights from formal systems and addresses limitations of current multi-agent architectures. Gödel’s second incompleteness theorem (Kennedy, 2008; Gödel, 1967) established that no consistent formal system can prove its own consistency. We extend this insight to LLMs, which operate on probabilistic rather than axiomatic foundations, making them inherently incapable of reliable self-validation. Table 1 illustrates how MACI’s distributed validation architecture addresses these fundamental limitations.
Current multi-agent systems (MAS) primarily serve as integration platforms for coordinating multiple LLMs. Notable frameworks include Microsoft’s AutoGen (Wu et al., 2024), the Multi-LLM Agent Debate Framework (Du et al., 2023; Chang, 2023, 2024a, 2024b), LangGraph and CrewAI (LangChain AI, 2024; Moura, 2024), XAgent (Xia et al., 2023), and CAMEL (Li et al., 2023). Although these frameworks excel at agent coordination, they focus on task distribution rather than comprehensive constraint management required for complex planning.
MACI’s meta-planning module (Section 4.1) addresses these limitations by acting as a metasystem that coordinates specialized agents. By decoupling planning and validation functions, MACI creates a validation framework analogous to the higher-order system required to demonstrate formal consistency while maintaining awareness of interdependent constraints.
3 Case Study: Can LLMs Handle Planning
This section begins by presenting various types of planning, including strategic approaches and tactical methodologies. We then present experimental results on a scheduling problem to illustrate how current LLMs fall short. Based on our diagnoses, we propose remedies in Section 4.
Planning Types
Planning can be classified into seven types. The first three are considered more strategic and the rest are tactic methodologies.
  1. 1.
    Sequential planning: Creates time-ordered schedules meeting goals/constraints via formal temporal representations (Allen & Hayes, 1989)
  2. 2.
    Proactive planning: Anticipates future scenarios to preemptively address issues through deliberate foresight (Cox & Veloso, 1998)
  3. 3.
    Retrospective planning: Improves future plans by analyzing past experiences via case-based reasoning (Kolodner, 1993)
  4. 4.
    Adaptive planning: Adjusts plans to changing conditions while preserving core constraints (Hammond, 1990)
  5. 5.
    Reactive planning: Prioritizes immediate actions over long-term goals in dynamic, uncertain environments (Georgeff & Lansky, 1987)
  6. 6.
    Predictive planning: Uses data-driven models to forecast scenarios and optimize decisions (Kushmerick et al., 1995)
Problem Statement
We conduct experiments using a Thanksgiving dinner planning problem designed as follows:
Initial Setup:
  • Mom (Sarah) is hosting Thanksgiving dinner at 6:00 PM in Boston. The following family members are traveling:
  • Dad (James) flying from San Francisco, landing at 1:00 PM Eastern time.
  • Sister (Emily) flying from Chicago, landing at 2:30 PM
  • Brother (Michael) driving from New York, estimated arrival 3:00 PM at home
  • Grandma is healthy and needs to be picked up from her home in suburban Boston
Critical Dependencies:
  • James needs to rent a car after landing
  • Emily doesn’t drive, needs pickup from airport, no other transportation options are allowed
  • Turkey needs 4 hours to cook, someone must be in the house once turkey is in oven
  • Side dishes require 2 hours of preparation
  • Travel time between home and BOS airport is one hour
  • Travel between BOS and grandma home is one hour
  • Travel between home and grandma home 30 minutes
Planning Question Set:
  1. 1.
    At what time should cooking start?
  2. 2.
    Who should pick up Emily and when?
  3. 3.
    When should Grandma be picked up by who?
Using this problem, three key types of planning conducted by GPT4o and DeepSeek are examined: sequential, proactive/reactive, and retrospective.
Table 2: DeepSeek Proposed Schedule
Time Task Assigned
12:00pm Sarah starts preparing side dishes. Sarah
1:00pm James arrives at the airport, begins car rental (takes 30-60 minutes). James
2:00pm James is ready to pick up Emily; Sarah and James put the turkey in the oven. J, S
2:30pm Emily lands at BOS. Emily
3:00pm James arrives at airport to pick up Emily J
3:00pm Michael arrives at home and leaves to pick up Grandma. Michael
3:15pm Michael at Grandma’s home. M
3:45pm Grandma arrives home; Michael returns. M
4:00pm James and Emily arrive home. J, E
6:00pm Dinner is served. All
3.1 Study #1 Sequential Planning
GPT-4o and DeepSeek struggled with real-world travel logistics, omitting key human considerations such as luggage claim time, rental car processes, and spatial relationships between airport terminals and the rental center (Table 2). These gaps forced manual constraint additions, highlighting LLM’s inability to integrate experiential knowledge, a gap addressed by the MACI common sense agent.
DeepSeek’s schedule further revealed spatial-temporal errors: 1) Spatial: Assumed James drove home immediately after renting a car at Boston Logan, ignoring his airport location while awaiting Emily; and 2) Temporal: Directed Michael to return home before heading to Grandma’s, missing the optimal direct route from NYC.
Table 3: GPT4o Proposed Schedule
Time Task Assigned
1:00pm James lands in Boston James
2:00pm Turkey goes into the oven Sarah
2:00pm James finishes car rental J
2:30pm Emily lands at BOS Emily
2:30pm James picks up Emily at airport J
3:00pm Michael arrives home Michael
4:00pm Side dishes preparation starts S, M
5:00pm Michael leaves to pick up Grandma M
5:30pm Michael arrives home with Grandma M
6:00pm Dinner is served All
Table 3 shows the GPT-4o schedule, which appears feasible but contains two critical errors in the case of adaptive planning required for emergency: 1) Arithmetic: Incorrectly calculates Grandma’s round-trip driving time as 30 minutes (vs. 30×2 minutes); and 2) Over-Constraint: Assumes only Sarah must watch the oven (vs. “someone”), creating brittleness under reduced slack time (e.g., delays).
Analysis (Appendix A) links both errors to flawed reasoning in constraint interpretation.
Diagnoses: Common-Sense Constraints and Isolated Processing Syndrome
Current LLM systems require explicit specification of real-world constraints that humans consider common sense, highlighting a limitation in their planning capabilities. Furthermore, we identified what we term isolated processing syndrome, where LLMs handle sub-tasks independently without maintaining awareness of overall constraints. This syndrome manifests itself in two critical ways: the system either misses obvious optimizations or proposes solutions that violate the stated constraints, leading to an infeasible or suboptimal plan.
3.2 Study #2 Proactive and Reactive Planning
Real-world scenarios do not always follow plans precisely. Robust systems require contingency planning for factors such as weather, traffic, or airline changes. These cascade through schedules, demanding adaptive replanning.
Table 4: GPT4o Revised Thanksgiving Schedule. Hazard! No one home watch oven between 3:00pm and 4:00pm.
Time Task Assigned
2:00pm Turkey placed in oven (4-hour cooking time begins) Sarah
3:00pm Michael arrives home Michael
Michael departs to pick up Emily from airport Michael
3:00pm Sarah departs to pick up Grandma Sarah
3:30pm Arrive at Grandma’s house Sarah
4:00pm Arrive at airport for Emily’s pickup Michael
Sarah home with Grandma -
James’s flight lands James
Begin side dish preparation Sarah
4:30pm James completes car rental process James
5:00pm Michael returns home with Emily -
5:30pm James arrive home -
6:00pm Thanksgiving dinner served Everyone
To evaluate LLMs’ dynamic replanning, we introduce a major disruption in our Thanksgiving scenario: James’s flight is delayed by 3 hours (arrival 4:00 PM vs. 1:00 PM). This forces adjustments to pickups, meal prep, and coordination while preserving original constraints.
LLM responses reveal critical flaws: 1) DeepSeek violates core constraints by unjustifiably delaying dinner to 7:00 PM (vs. the 6:00 PM deadline); and 2) GPT-4o (Table 4) commits a safety violation: leaving the oven unattended, despite explicit constraints. These errors highlight LLMs’ inability to reliably maintain and validate constraints during replanning, even with full information.
Diagnosis: Attention Narrowing
Claude detects constraint violations in other LLMs’ plans but both GPTo4 and DeepSeek struggle with self-validation, revealing an asymmetry in error detection. LLMs often embed flawed interpretations of constraints during planning (e.g., rigidly interpreting “someone must be in the house” when the turkey is in the oven), propagating errors through their frameworks.
Two key limitations emerge: 1) Attention narrowing: Overfocus on objectives (e.g., arrival times) causes neglect of critical constraints (e.g., fire safety); and 2) Solution rigidity: Once a constraint is satisfied (e.g., assigning Sarah to oven duty), LLMs treat it as a fixed context, failing to explore alternatives.
More specifically, GPT-4o assigned Sarah to monitor the oven, but missed reallocating this task to Grandma earlier, preventing Sarah from serving as an additional driver, a missed efficiency gain.
3.3 Study #3 Retrospective Planning
Retrospective planning improves future strategies by learning from past decisions. Both DeepSeek and GPT-4o lacked flexibility even after error correction: DeepSeek relied on rigid templates, and GPT-4o’s initial assignment (Sarah homebound in Table 3) limited later options.
Missed Opportunities During James’ 3-Hour Delay:
  • Sarah could have picked up Grandma earlier, freeing her for contingencies like Emily’s transport.
  • Michael could have left NYC at 10:00 AM to reach Emily by 2:30 PM.
In Section 4, we show how MACI’s meta-planner effectively revises plans by debugging errors and adapting to dynamic constraints.
3.4 Summary of LLM Limitations in Planning
Our analysis reveals three core limitations in current LLMs and reasoning methods (CoT (Wei et al., 2022), ToT (Yao et al., 2023)):
Metacognitive Limitations
LLMs struggle with self-validation and constraint awareness. While external LLMs detect errors in others’ plans, planners consistently miss their own violations (e.g., GPT-4o rigidly assigned Sarah to oven duty without considering Grandma’s availability). Key causes are:
  1. 1.
    Pattern-matching optimization vs analytical validation
  2. 2.
    No belief-state tracking during reasoning
  3. 3.
    Single-solution focus vs comparative analysis
Current reasoning methods exacerbate these issues by operating within the same flawed cognitive framework.
Attention Bias
Transformer architectures prioritize recent context, creating: 1) Narrowing: Recent constraints (arrival times) overshadow earlier ones (oven safety); and 2) Isolated Processing: Sub-tasks addressed without holistic awareness
Common Sense Gaps
LLMs miss implicit real-world knowledge (luggage claim times, rental car logistics) requiring explicit specification of human-obvious constraints (airport-terminal proximity).
4 MACI Framework Specification
MACI implements a three-layer architecture to address current LLM limitations: metacognitive constraints, attention bias, and gaps in common-sense reasoning. The top two layers represent key innovations, while the bottom layer leverages existing multi-agent system (MAS) infrastructure.
4.1 Three-Layer Architecture
Meta-Planner: The Planning Generator
The meta-planner $\mathcal{MP}$ functions as a higher-order planner that generates task-specific planning systems:
$\mathcal{MP}:(\mathcal{O},\mathcal{C_{E}})\rightarrow\mathbf{W},$
where $\mathbf{W}$ represents a planning system composed of specialized, coordinated agents. Similar to how a compiler generator produces compilers from specifications, $\mathcal{MP}$ constructs agent networks from task requirements. It analyzes objectives, identifies required roles and dependencies, selects appropriate agents, and establishes interaction protocols. This produces a complete workflow template defining both the planning state space and the agent coordination mechanisms needed to solve the given task.
Agent Repository Layer
This layer maintains a distributed collection of planning agents, each designed with a restricted context window and specialized interface. By dividing cognitive tasks among agents, this architecture ensures a complete knowledge of the constraints without overwhelming individual components.
System Infrastructure Layer
Built on open-source MAS frameworks, the infrastructure layer supports essential operations such as agent registration, message routing, resource allocation, and deployment scaling.
4.2 Agent Repository Layer Design
4.2.1 Agent Registration and Search Protocol
Agents register in the repository with five attributes: input/output protocol ($P$), type ($t$), capability vector ($c$), context window size ($w\leq 1k$ tokens) and user rating ($r$). These attributes enable $\mathcal{MP}$ to match task requirements to agent capabilities efficiently. $\mathcal{MP}$ searches for agents by: 1) matching task requirements to capability descriptions ($c$), 2) filtering by agent type ($t$) and user ratings ($r$), and 3) verifying protocol compatibility ($P$).
4.2.2 State Space and Agent Design
General State Space
Table 5: State Space Dimensions and Components
Dimension Core Components Example States
Who
• Actor ID
• Current Roles
• Role Constraints
• Role Changes
• Driver vs Passenger
• Supervisor vs Worker
• Buyer vs Seller
• Multiple role conflicts
Where
• Current Position
• Target Location
• Transition Points
• Access Rules
• Physical locations
• Virtual positions
• State transitions
• Boundary constraints
When
• Time Points
• Duration
• Deadlines
• Probability
• Event timestamps
• Process duration
• Completion times
• Delay likelihood
What
• Methods
• Requirements
• Constraints
• Costs
• Tools/Vehicles
• Tickets/Permits
• Capacity limits
• Time-money trade-offs
Why
• Rationale
• Dependencies
• Risks
• Alternatives
• Decision basis
• Causal chains
• Failure modes
• Backup plans
Table 5 presents five dimensions of state space, Who, Where, When, What, and Why, that form the basis of planning tasks. These dimensions are general enough to apply across domains, with travel planning serving as an illustrative example.
Common Agents
Common agents are designed to handle core planning and validation tasks that are universally applicable across domains. These agents provide foundational capabilities that ensure the consistency, feasibility, and robustness of planning systems, regardless of the specific application. Their roles align with the general state space dimensions of Who, Where, When, What, and Why, ensuring that all essential aspects of planning are addressed. For a list of common agents, please refer to Appendix D.2.
These agents work collaboratively to ensure robust planning and validation across diverse domains. Although the list of common agents is not exhaustive, it provides a strong foundation to address most planning tasks. New common agents can be introduced as the need arises, provided they offer capabilities that can be reused across multiple domains. For example, an Adaptation Agent could be added to dynamically adjust constraints in response to evolving task environments.
Specialized Agents
Specialized agents extend the capabilities of MACI by addressing domain-specific requirements that are not generalizable across all tasks. These agents are designed to implement detailed knowledge and expertise tailored to specific applications, ensuring that MACI remains adaptable and effective for diverse planning challenges. Specialized agents typically focus on tasks requiring unique constraint handling, optimization, or analysis.
Examples of specialized agents include the following.
  • Travel Planning Agent: Optimizes itineraries, manages transportation options, and ensures seamless transitions between locations.
  • Financial Risk Agent: Monitors market trends, assesses risks, and validates investment strategies.
  • Healthcare Resource Agent: Allocates medical supplies, schedules staff, and optimizes patient care pathways.
  • Supply Chain Agent: Manages logistics, tracks inventory, and optimizes distribution routes.
Specialized agents adhere to architectural principles that maintain consistency and integration within MACI:
  • Restricted Context Windows: Each agent operates within a limited context to ensure efficiency and focus ($w\leq 1k$ tokens).
  • Protocol Buffer Interfaces: Standardized interfaces enable seamless communication between agents.
  • Coordination with Common Agents: Specialized agents leverage the capabilities of common agents for validation, compliance, and constraint management.
  • Domain-Specific Constraints: Specialized agents are equipped to handle unique requirements and challenges specific to their domain.
4.2.3 Agent Coordination and State & Constraint Management
MACI ensures seamless coordination between agents across state dimensions. For example, in travel planning, when a Schedule Agent detects a flight connection constraint, it triggers updates to Location, Resource, and Role Manager Agents. These updates propagate through a state transition protocol, where the proposed state change is validated against global and local constraints:
$T:S_{t}\rightarrow S_{t+1}\text{ subject to }C_{\text{global}}\land C_{\text{local}}$
4.2.4 Validation Protocols
The validation agent ensures state consistency by decomposing global constraints into dimension-specific checks. For example, when booking a flight, the temporal agent validates the temporal constraint, the spatial feasibility is checked by the location agent, and the availability of seats is validated by the resource agent. Common sense agents identify overlooked constraints, such as luggage claiming time, inter-terminal walking time, and rental car pickup time. If validation fails, the system triggers a recovery process to generate alternative solutions. Validation protocols can be updated when unexpected events occur to improve retrospective planning.
4.2.5 Other Considerations
Due to space limitation, we document in Appendix C additional considerations including cross-domain generalization, enhancement and integration of common sense knowledge, scalability, and a list of future enhancements when more applications will be supported.
4.3 Meta-Planner: Planning a Planner to Plan
The meta-planner $\mathcal{MP}$ operates as a higher-level planner that constructs other planners, which are actionable workflows for specific tasks.
4.3.1 The Meta-Planner Algorithm
$\mathcal{MP}$ is specified in Algorithm #1. A list of all symbols used in the specifications is provided in Table 20 in Appendix D.
A workflow network $\mathbf{W}=(\mathcal{N},\mathcal{E})$, represented as a directed graph, consists of roles as nodes and their dependencies as edges. The nodes $n\in\mathcal{N}$ denote the roles performed by human actors (e.g. driver, cook, supervisor), while the edges $e_{ij}\in\mathcal{E}$ between the nodes $n_{i}$ and $n_{j}$ capture the monitored dependencies (e.g., temporal, spatial constraints, data dependencies and safety requirements for supervision).
$\mathbf{W}=(\mathcal{N},\mathcal{E}),\text{ where }\mathcal{N}=A^{*}_{n}\text{and }\mathcal{E}=A^{*}_{e}.\vspace{-.1in}$ (1)
Algorithm 1 $\mathcal{MP}$: Planner for Planning a Plan
0:  Objectives $\mathcal{O}$, explicit constraints $\mathcal{C_{E}}$, agent pool $\mathbf{A}$, people $\mathcal{P}$, metrics $\mathcal{M}$
0:  Optimized workflow $\mathbf{W^{*}}=(\mathcal{N},\mathcal{E})$ (Eq. 1)
  // Phase 1: Network Construction
  1. $\mathcal{N}$: Extract roles & capabilities from $\mathcal{O}$ (Eq. 2)
  2. $\mathcal{E}$: Identify node dependencies from $\mathcal{C_{E}}$ (Eq. 3)
  // Phase 2: Agent Assignment
  3. Assign role agents: $\forall n\in\mathcal{N}$, select $\alpha_{n}\in\mathbf{A_{n}}$ (Eq. 4)
  4. Assign edge agents: $\forall e_{ij}\in\mathcal{E}$, $\alpha_{ij}\in\mathbf{A_{e}}$ (Eq. 5)
  // Phase 3: Iterative Refinement
  while improvement in $V(\mathbf{W},\mathcal{M})$ do
     // Role Assignment Optimization
     for all $n\in\mathcal{N}$ do
        Update role-person mappings via $f_{\text{role}}(n,\mathcal{P})$
     end for
     // Dependency Monitoring
     for all $e\in\mathcal{E}$ do
        Verify dependencies via assigned edge agents
     end for
     // Validation
     if $V(\mathbf{W_{new}},\mathcal{M})>V(\mathbf{W_{current}},\mathcal{M})$ then
        $\mathbf{W_{current}}\leftarrow\mathbf{W_{new}}$
     end if
  end while
  return $\mathbf{W^{*}}=\mathbf{W{current}}$
$\mathcal{MP}$ constructs a planner in three phases. First, it builds the network structure by identifying roles and their dependencies, forming the foundation of the workflow.
Second, it assigns appropriate agents to manage both roles and dependencies. Each node $n\in\mathcal{N}$ is managed by a designated role agent $\alpha_{n}$. Similarly, each edge $e_{ij}\in\mathcal{E}$, representing a dependency between nodes, is managed by an agent $\alpha_{ij}$, which connects the node agents $\alpha_{i}$ and $\alpha_{j}$. These edges are enriched with metadata that may include details such as data types, communication protocols, timing constraints, and security requirements.
Finally, the generated planner iteratively refines the workflow. This process involves optimizing role assignments and continuously monitoring dependencies until the validation score stabilizes, ensuring a well-validated workflow.
4.3.2 Meta-Planning Components
Let us clearly distinguish between roles (requiring human qualifications) and monitoring tasks (requiring capabilities).
Role and Qualification Analysis
$\mathcal{MP}$ analyzes objectives $\mathcal{O}$ to identify required roles and their qualifications:
$\text{map}_{\text{role}}:\mathcal{O}\rightarrow\{(n_{i},q_{i})\}$ (2)
where $n_{i}$ represents a role and $q_{i}$ its required qualifications. For example, $(n_{\text{driver}},q_{\text{license}})$ or $(n_{\text{cook}},q_{\text{experience}})$.
Constraint Management
The framework maintains a global constraint set $C$ composed of:
$C=C_{E}\cup C_{I}\cup C_{D}$ (3)
where $C_{E}$ represents explicit constraints from problem specification, $C_{I}$ denotes implicit constraints identified by Common Sense Agent, and $C_{D}$ represents derived constraints from agent interactions.
Agent Assignment
Two types of agents, node and edge, are selected from the repository $\mathbf{A}$:
1. Node Monitoring Agents:
$A^{*}_{n}=\operatorname*{arg\,min}_{A_{i}\in\mathbf{A}}\sum_{n_{j}}\text{dist}(q_{j},A_{i}.\text{capabilities})$ (4)
These agents verify role qualifications and manage people-role assignments and transitions.
2. Edge Monitoring Agents:
$A^{*}_{e}=\operatorname*{arg\,min}_{A_{i}\in\mathbf{A}}\sum_{e_{j}}\text{dist}(c_{j},A_{i}.\text{capabilities})$ (5)
These agents monitor dependencies between roles, where $c_{j}$ represents constraint types (temporal, spatial, etc.).
$\mathcal{MP}$ iteratively refines the network by validating workflow constraints and performance and replacing or adding agents, i.e., checking and adding nodes and edges.
4.4 Workflow Execution Framework
The workflow $\mathbf{W^{*}}$ requires a runtime system for execution. In this work, we evaluate $\mathbf{W^{*}}$ by entering $\mathbf{W^{*}}$ along with the problem statement into an LLM (DeepSeek and GPT4o). A key limitation is that the feedback loop for refining $\mathbf{W^{*}}$ must be performed manually.
5 Evaluating $\mathcal{MP}$’s Planning Capabilities
To assess $\mathcal{MP}$’s performance and adaptability, we adopted a dual-approach experimental structure. The first experiment uses the Traveling Salesperson Problem (TSP) to validate $\mathcal{MP}$’s optimization capabilities. The second involves the Thanksgiving Dinner Planning problem, showcasing $\mathcal{MP}$’s ability to handle complex, real-world challenges with cross-thread dependencies and dynamic adaptability.
Traveling Salesperson Problem (TSP)
The TSP experiment benchmarks $\mathcal{MP}$ against planners (Claude, DeepSeek, GPT4o) and their $\mathcal{MP}$ integrations. Metrics include: 1) solving quality, and 2) solution optimality. Detailed results are in Appendix E. The conclusion is that though the task is simple, $\mathcal{MP}$ helps improve results for monitoring execution, validating results, and suggesting algorithmic approaches.
Thanksgiving Dinner Planning
This problem, detailed in Section 3, evaluates $\mathcal{MP}$’s ability to generate workflows $\mathbf{W^{*}}$ with enhanced constraints and dependencies. It involves multiple agents performing interdependent tasks, representing a complex multi-agent scenario.
Planning performance is compared across three configurations: DeepSeek + $\mathcal{MP}$, GPT4o + $\mathcal{MP}$, and Claude + $\mathcal{MP}$. Metrics include:
[…]
where flexibility measures slack time incorporated to handle unexpected events.
5.1 Meta-Planning for Thanksgiving Event
Following Algorithm 1, $\mathcal{MP}$ generates workflows with:
  • Role nodes (e.g., cook, drivers, supervisor),
  • Explicit constraint edges (e.g., temporal, spatial, etc.),
  • Implicit constraint edges from common-sense analysis.
The planner monitors nodes and edges, enabling dynamic adjustments. Full specifications are in Appendix A.
Evaluation Scenarios
We test $\mathcal{MP}$ under:
  1. 1.
    Sequential Planning: Task executed as planned.
  2. 2.
    Reactive Planning: A 3-hour flight delay requiring task reallocations.
Meta-Planner Output
$\mathcal{MP}$ enhances planning by:
  • Identifying implicit constraints (e.g., luggage claim time, car rental delays),
  • Clarifying role dependencies,
  • Incorporating common-sense constraints (e.g., fatigue, social preferences),
In reactive planning, $\mathcal{MP}$ integrates an early information agent to detect flight delays at departure, enabling timely workflow updates and demonstrating adaptability.
5.2 Experimental Results
Sequential Planning Performance
With $\mathcal{MP}$’s enhanced workflow $\mathbf{W^{*}}$, all three LLMs successfully generated feasible solutions, a significant improvement over their previous failures with the original problem specification.
As illustrated in Table 6, DeepSeek demonstrated superior scheduling efficiency by optimizing James’s airport wait time for Emily’s pickup, requiring only two iterations. While GPT4o eventually produced a valid solution in three iterations, it created suboptimal travel patterns by having Michael make separate trips. Claude’s solution, though feasible in two iterations, included unnecessary travel between pickup tasks. This experiment highlighted how $\mathcal{MP}$’s explicit constraint specification and common-sense augmentation enabled consistent performance improvement across different LLMs.
Table 6: Sequential Planning Performance. (# = iterations)
LLM # Notable Features
DeepSeek 2 Optimized airport wait time for James; balanced workload
GPT4o 3 Extra travel for Michael; suboptimal load balance
Claude 2 Unnecessary travel between pickup tasks
Reactive Planning Performance
GPT4o failed to produce a valid solution, violating three critical constraints without the ability to recognize these violations.
Reactive Planning Performance
The flight delay scenario revealed significant differences between LLMs’ capabilities. DeepSeek demonstrated superior spatial reasoning by routing Michael directly to the airport, an insight that should have come from $\mathcal{MP}$’s common-sense spatial reasoning. This unexpected ability to improve workflow highlights the synergy between $\mathcal{MP}$ and LLM —$\mathcal{MP}$ provided early alert through its information agent (Table 15 in Appendix A.9), which DeepSeek leveraged for immediate replanning.
In contrast, Claude produced two feasible plans but missed the 10:00 AM alert in $\mathbf{W^{R}}$, starting its schedule at 1:00 PM and missing opportunities for proactive actions like early Grandma pickup to free Sarah as a driver. GPT4o failed entirely, producing three constraint violations it couldn’t recognize, preventing further improvements.
Table 7: Reactive Planning Performance (Flight Delay)
LLM # Notable Features
DeepSeek 3 Smart routing of Michael directly to airport; efficient travel patterns
GPT4o Failed Failed to maintain critical constraints; unable to recover
Claude 3 Two valid plans with different trade-offs; longer wait times
6 Conclusions with Impact Statements
Current artificial intelligence systems excel at pattern matching but struggle with deliberate analytical reasoning required for complex planning. While Large Language Models (LLMs) demonstrate impressive capabilities, they face three critical limitations: inability to validate their own outputs (paralleling Gödel’s incompleteness theorems), inconsistent maintenance of constraints during planning, and inadequate integration of common-sense knowledge. Our experiments demonstrate that existing multi-agent systems, focused primarily on coordination, fail to address these limitations.
MACI (Multi-Agent Collaborative Intelligence) advances the state of the art through three key contributions:
  1. 1.
    Meta-Planning Architecture: MACI generates specialized planning workflows by:
    • Constructing role-node and dependency networks.
    • Identifying implicit constraints through common-sense analysis.
    • Establishing validation across distributed agents.
  2. 2.
    Workflow Enhancement: As demonstrated in our experiments:
    • Sequential Planning: Enhanced workflow $\mathbf{W^{*}}$ enabled all LLMs to generate feasible solutions.
    • Reactive Planning: Early information agent in $\mathbf{W^{R}}$ enabled proactive replanning.
    • Unexpected synergy: DeepSeek contributed spatial optimizations beyond MACI’s specifications.
  3. 3.
    Empirical Validation: Our Thanksgiving dinner planning case study shows:
    • Improving constraint satisfaction on LLMs.
    • Effective handling of crises through early detection.
    • Clear performance differences between LLMs in utilizing MACI’s capabilities.
MACI demonstrates that addressing LLM limitations requires more than simple agent coordination. It demands a structured meta-planning architecture that enables comprehensive constraint awareness and validation. The framework’s success in complex temporal planning, coupled with unexpected improvements from LLM contributions, suggests a promising direction for developing more capable AI planning systems.
Appendices
Appendix A Experiment Details: Meta-Planning for the Thanksgiving Dinner Task
The problem statement remains consistent with Section 3, with $\mathbf{W^{*}}$ generated by $\mathcal{MP}$ to enhance constraints and dependencies. Planning performance is compared across four configurations: DeepSeek, GPT4o, DeepSeek + $\mathcal{MP}$, and GPT4o + $\mathcal{MP}$.
A.1 Phase 1: Network Construction
A.1.1 Node (Role) Specifications
First, meta-planner $\mathcal{MP}$ extracts roles ($\mathcal{N}$) with their required qualifications:
  • $n_{\text{cook}}$: capability to prepare dinner
  • $n_{\text{driver1}}$: capability to drive, pick up from airport
  • $n_{\text{driver2}}$: capability to drive, pick up grandma
  • $n_{\text{supervisor}}$: capability to monitor oven
A.1.2 Edge (Dependency) Specifications
Next, $\mathcal{MP}$ identifies dependencies ($\mathcal{E}$) between roles:
=$\mathcal{E}=\{e_{\text{temporal}},e_{\text{spatial}},e_{\text{safety}}\}$ (6)
The critical dependencies include:
  • $e_{\text{temporal}}$: - Turkey (4 hours) must finish by 6:00 PM - Side dishes (2 hours) must finish by 6:00 PM - Airport pickups must align with landing times
  • $e_{\text{spatial}}$: - Driver-passenger location matching - Travel time constraints between locations
  • $e_{\text{safety}}$: - Continuous oven supervision requirement
A.2 Phase 2: Agent Assignments
After constructing the network structure, $\mathcal{MP}$ selects and assigns agents to monitor both the roles and dependencies.
A.2.1 Node (Role) Agent Assignment
For each role, $\mathcal{MP}$ selects monitoring agents with the required capabilities:
$f_{\text{role}}:\mathcal{N}\rightarrow\mathbf{A}$ (7)
The role monitoring agents include:
  • Cook Monitor: Tracks cooking timeline, coordinates meal components
  • Driver Monitor: Validates driver availability
  • Supervisor Monitor: Ensures oven supervision
  • Resource Monitor: Manages vehicle assignments and actor schedules
A.2.2 Edge (Dependency) Agent Assignment
For the identified dependencies, $\mathcal{MP}$ assigns specialized monitoring agents:
$f_{\text{edge}}:\mathcal{E}\rightarrow\mathbf{A}$ (8)
Dependencies require these monitoring agents:
  • Temporal Agent: Manages timing constraints (cooking durations, travel times, arrival schedules)
  • Spatial Agent: Tracks location constraints (airport-home-grandma routes)
  • Safety Agent: Ensures oven supervision constraint remains satisfied
The resulting agent assignments create a complete monitoring system where:
  • Role agents track individual actor assignments and qualifications
  • Edge agents monitor interactions and dependencies between roles
  • All agents coordinate to maintain global constraint satisfaction
Table 8: Node (Role) Monitoring Agent Requirements
Agent Input Protocol Output Protocol
Cook
Monitor
Role: cook
Qualifications: skills
Time: prep and cook
Status: progress
Alerts: timing issues!
Updates: completed?
Driver
Monitor
Role: driver
Qs: license, rest
Where: current GPS
Status: availability
Alerts: fatigue warnings
Updates: new GPS
Supervisor
Monitor
Role: supervisor
Location: house
Duration: cover time
Status: covered?
Alerts: coverage gaps!
Updates: role transitions
Table 9: Edge (Dependency) Monitoring Agent Requirements
Agent Input Protocol Output Protocol
Temporal Start times
Durations
Deadlines
Buffer requirements
Schedule conflicts
Timing violations
Schedule updates
Spatial Locations
Routes
Travel times
Traffic conditions
Route violations
Location conflicts
Path updates
Safety Critical constraints
Resource states
Coverage requirements
Safety violations
Resource conflicts
Mitigation plans
A.2.3 Common Sense Constraint Analysis (Performed by an LLM)
A common sense agent identifies the following implicit constraints that can affect Thanksgiving dinner planning. This list is generated by Claude given the problem statement.
  • Physical Processing Times:
    • Airport luggage claim: 30 minutes
    • Car rental procedures: 30 minutes
    • Holiday traffic variations
    • Winter weather considerations
  • Human Factors:
    • Driver fatigue after long trips
    • Cooking preparation overhead
    • Multi-tasking limitations
    • Task switching delays
    • Required rest periods
  • Resource Dependencies:
    • Vehicle passenger capacity
    • Oven temperature management
    • Kitchen workspace limits
    • Shared resource coordination
  • Social Considerations:
    • Personal preferences for interactions
    • Family dynamics in assignments
    • Post-travel guest comfort
    • Host preparation requirements
Table 10: Complete Workflow Specification: Nodes, Edges, and Agent Assignments
Type Component Requirements Agent Protocol Dependencies
Node Components (Roles)
Node Cook Role
- Turkey (4hr)
- Side dishes (2hr)
- Kitchen management
- Time management
Input: schedule, resources, recipes
Output: task progress, completion
Monitor: kitchen_state() → status
Validate: cooking_constraints()
Connected to:
- Supervisor
- Resource edges
Node Driver1
(James/Michael)
- Valid license
- Airport navigation
- Car rental capable
- Rest state adequate
Input: flight times, routes
Output: location, ETA
Monitor: driver_state() → status
Validate: driver_constraints()
Connected to:
- Airport pickup
- Travel edges
Node Driver2
- Valid license
- Local navigation
- Availability window
- Rest state adequate
Input: pickup schedule, route
Output: location, ETA
Monitor: driver_state() → status
Validate: driver_constraints()
Connected to:
- Grandma pickup
- Travel edges
Node Supervisor
- Home presence
- Oven monitoring
- Safety awareness
- Time commitment
Input: cooking schedule, rules
Output: supervision status
Monitor: safety_state() → status
Validate: safety_constraints()
Connected to:
- Cook role
- Safety edges
Edge Components (Dependencies)
Edge Temporal - Schedule tracking
- Buffer management
- Sequence logic
- Critical path
Input: timestamps, durations
Output: schedule conflicts
Monitor: schedule_state() → alerts
Optimize: timeline_adjust()
Connects:
- All roles
- All activities
Edge Spatial - Location tracking
- Route optimization
- Traffic updates
- Distance constraints
Input: locations, routes
Output: travel updates
Monitor: location_state() → alerts
Optimize: route_adjust()
Connects:
- Drivers
- Locations
Edge Resource - Vehicle allocation
- Kitchen resources
- People availability
- Capacity limits
Input: resource demands
Output: allocation status
Monitor: resource_state() → alerts
Optimize: resource_adjust()
Connects:
- All roles
- All resources
Edge Safety - Oven monitoring
- Driving safety
- Food safety
- Critical rules
Input: safety requirements
Output: violation alerts
Monitor: safety_state() → alerts
Enforce: safety_rules()
Connects:
- All roles
- Critical tasks
A.2.4 Common Sense Constraint Analysis and Verification (Human in the Loop)
The common sense constraints identified above require different verification approaches:
Agent-Required Information
These constraints need specialized agents to verify and quantify:
  • Airport Operations
    • United Airlines’ average luggage delivery time at BOS Terminal B
    • Terminal B to rental car center: shuttle schedule, walking options
    • Historical flight delay patterns for November at BOS
  • Weather and Traffic
    • Boston weather forecast for the event date
    • Historical traffic patterns on Thanksgiving days
    • Impact on airport-city-suburb travel times
  • Task Dependencies
    • Kitchen workflow analysis for parallel cooking tasks
    • Resource contention in meal preparation
    • Critical path identification in cooking timeline
Human Verification
Certain constraints require explicit human input to ensure that the planning process takes into account subtle interpersonal and individual considerations. These include:
  • Family Dynamics
    • Preferred pickup arrangements for Grandma (e.g., Grandma loves to have a grandson surprise her).
    • Optimal relationship-based task pairings.
    • Social comfort factors in assignments (e.g., Sarah and Grandma do not work together in the kitchen).
  • Personal Capabilities
    • Individual cooking experience levels.
    • Driver comfort with airport navigation.
    • Multi-tasking abilities of participants.
This separation ensures that agents focus on collecting quantifiable data while humans provide essential social and personal insights. $\mathcal{MP}$ can then integrate both types of information into the final workflow design.
A.3 Agent Requirements and Assignments
The $\mathcal{MP}$ requires two categories of agents. $\mathcal{MP}$ specifies their requirements in the protocol buffer format in Table 8 for the nodes and Table 9 for the edges, respectively.
Each agent must implement these protocols to participate in the workflow. The meta-planner selects agents from the pool based on their ability to satisfy these interface requirements. During execution, agents communicate through these standardized protocols while maintaining their specialized monitoring functions.
A.4 Monitoring Protocols and Dynamic Adjustments
The workflow monitoring operates through a hierarchical protocol system that enables both routine supervision and dynamic adjustments.
Basic Monitoring Protocol
Each agent maintains a continuous monitoring cycle:
[…]
For example, the temporal agent tracks schedule adherence:
$\Delta t=t_{\text{planned}}-t_{\text{actual}}\begin{cases}\text{normal}&\text{if }|\Delta t|<\text{buffer}\\ \text{warning}&\text{if }\text{buffer}\leq|\Delta t|<\tau\\ \text{violation}&\text{if }|\Delta t|\geq\text{ threshold }\tau\end{cases}$ (10)
Dynamic Adjustment Mechanism
When deviations occur, the system initiates a three-phase response:
1. Impact Assessment:
$\text{impact}(e)=\sum_{n\in\text{affected}(e)}\text{severity}(n)\times\text{urgency}(n)$ (11)
2. Solution Generation:
$S^{*}=\operatorname*{arg\,min}_{s\in\text{Solutions}}\{\text{cost}(s)|\text{feasible}(s)\}$ (12)
3. Coordination Protocol:
$\text{update}:(W_{\text{current}},S^{*})\rightarrow W_{\text{new}}$ (13)
For instance, if James’s flight is delayed:
  • Spatial agent detects arrival time change
  • Temporal agent calculates ripple effects
  • Role agents evaluate reassignment options
  • Safety agent verifies continued supervision coverage
The meta-planner $\mathcal{MP}$ coordinates these responses while maintaining global constraint satisfaction.
A.5 Integrated Workflow Network
Table 10 presents the resulting workflow network $\mathbf{W^{*}}$, which includes all nodes and edges, and their assigned agents and protocols.
  1. 1.
    Role Nodes:
    • Cook1: Sarah (primary) or Grandma (if at home) with 4-hour turkey + 2-hour sides
    • Driver1: James (after car rental) or Michael
    • Driver2: Available person after initial pickups
    • Supervisor: Must be present while turkey cooks
  2. 2.
    Dependencies:
    • Temporal: Verified airport processing + travel times
    • Spatial: Traveling routes with traffic consideration
    • Safety: Continuous oven supervision requirement
  3. 3.
    Agent Monitoring:
    • Temporal Agent: Schedules with verified buffer times
    • Spatial Agent: Real-time location and route mgmt.
    • Safety Agent: Role coverage for supervision
A.6 Agent Interaction Specifications
Please, see Table 11.
Table 11: Agent Interaction Protocols and State Transitions
Interaction Type Protocol State Transitions Validation Rules
Node-to-Node Interactions
[…]
Message: (task, duration, reqs.)
States: prep → cooking → comp.
Trigger: task_state_change()
Validate: coverage()
Alert: coverage_gap()
Driver1 $\leftrightarrow$ Driver2 Protocol: pickup_handoff()
Message: (location, time, passenger)
States: available → enroute → comp.
Trigger: location_change()
Validate: timing_feasible()
Alert: schedule_conflict()
Edge Agent Operations
Temporal Agent Protocol: schedule_monitor()
Message: (event, time, dependencies)
States: scheduled → active → comp.
Trigger: time_milestone()
Validate: timing_feasible()
Alert: delay_impact()
Spatial Agent Protocol: location_track()
Message: (actor, position, dest.)
States: idle → moving → arrived
Trigger: position_update()
Validate: route_feasible()
Alert: travel_delay()
A.7 New Problem Statement Revised with $\mathbf{W^{*}}$
Given the $\mathbf{W^{*}}$ generated by MACI’s meta-planner $\mathcal{MP}$, the Thanksgiving Dinner Planning problem statement stated at the beginning of this section is revised as follows:
Initial Setup:
  • Mom (Sarah) is hosting Thanksgiving dinner at 6:00 PM in Boston. The following family members are traveling:
  • Dad (James) flying from San Francisco, landing at 1:00 PM Eastern time.
  • Sister (Emily) flying from Chicago, landing at 2:30 PM
  • Brother (Michael) driving from New York, estimated arrival 3:00 PM at home
  • Grandma is healthy and needs to be picked up from her home in suburban Boston
Critical Dependencies:
  • James must rent a car after landing
  • Emily must be picked up from airport, no other transportation options are allowed
  • Turkey needs 4 hours to cook, someone must be in the house once turkey is in oven for safety
  • Side dishes require 2 hours of preparation, which can overlap with turkey
  • Travel time between home and Boston airport is one hour (one-way)
  • Travel between Boston airport and grandma home is one hour (one-way)
  • Travel between home and grandma home 30 minutes (one-way)
* New Dependencies:
  • The airport luggage pickup time after landing is 30 minutes.
  • Renting a car takes 30 minutes.
  • One person can simultaneously prepare turkey and side dishes.
  • Grandma prefers Michael to pick her up, provided that it does not cause the dinner time delay.
  • Grandma and Sarah prefer not to cook together in the kitchen.
  • Traffic congestion is not factored into current planning.
Planning Question Set:
  1. 1.
    All tasks and dependencies must be strictly observed in the plan, or the plan fails.
  2. 2.
    Dinner time is strictly at 6:00 PM, all tasks must be completed by then (redundancy).
  3. 3.
    Account for the idle time of each person.
  4. 4.
    The schedule consists of three columns: time, task, and assigned person(s).
A.8 Experiment #1: Sequential Planner
Once after the original plan was revised by $\mathcal{MP}$ to include more specific details, clarify ambiguous explicit constraints, and define implicit constraints, the performance of the three LLMs used in the experiment improved significantly. When the augmented plan $\mathbf{W^{*}}$ was input into DeepSeek, GPT4o, and Claude, each model successfully generated a feasible plan within two to three iterations. (The case study in Section 3 shows that DeepSeek was confusing and GPT4o repeatedly committed constraint violations.)
Table 12: DeepSeek’s Plan, Two Iterations
Time Task Assigned
1:00 PM James lands at Boston James
1:00–1:30 PM James picks up luggage James
1:30–2:00 PM James rents a car James
2:00 PM Turkey in oven (4 hours; requires monitoring) Sarah
2:00–3:00 PM James waits at airport (idle) James
2:30 PM Emily lands at Boston Emily
2:30–3:00 PM Emily waits for luggage Emily
3:00 PM James picks up Emily James
3:00 PM Michael arrives home Michael
3:00 PM Michael departs to Grandma Michael
3:30 PM Michael picks up Grandma Michael
3:30–4:00 PM Michael drives back home with Grandma Michael
3:00–4:00 PM James drives Emily home (airport to home: 1 hour) James
4:00 PM James and Emily home James
4:00 PM M. and Grandma home Michael
4:00–6:00 PM Sarah prepares side dishes Sarah
6:00 PM Thanksgiving dinner begins All
A.8.1 Results: DeepSeek Wins
Upon closer examination of the number of iterations required to produce a feasible plan, DeepSeek and Claude each required one revision (two iterations), while GPT4o required two revisions (three iterations). In terms of scheduling quality, measured by slack time, total driving distance, and load balance, DeepSeek (Table 12) outperformed both Claude (Table 14) and GPT4o (Table 13). DeepSeek optimized time and effort by scheduling James to wait at the airport for 30 minutes to pick up Emily. In contrast, Claude scheduled James to drive home and then return to the airport to pick up Emily, resulting in unnecessary travel. GPT4o assigned James to return home and scheduled Michael to first pick up Emily and then proceed to pick up Grandma, leading to a less balanced load. A better solution to reduce travel time would have been to schedule Michael to pick up Emily first and then drive with her to Grandma’s home to pick up Grandma, allowing all three to return home together. This adjustment would save 30 minutes of driving time and improve Grandma’s overall happiness to see both grandchildren.
Table 13: GPT4o’s Plan, Three Iterations
Time Task Assigned
1:00 PM Land at BOS Airport James
1:00-1:30 PM Luggage pickup James
1:30-2:00 PM Rent car James
2:00 PM Start turkey Sarah
2:00-3:00 PM Drive home James
2:30 PM Land at BOS Airport Emily
3:00 PM Arrive home Michael
3:00-4:00 PM Drive to airport, pick up Emily Michael
4:00-5:00 PM Return home with Emily Michael
5:00-5:30 PM Drive to Grandma’s Michael
5:30-6:00 PM Return with Grandma Michael
4:00-6:00 PM Prepare side dishes Sarah
6:00 PM Dinner served All
Table 14: Claude’s Plan, Two Iterations
Time Task Assigned
1:00 PM Land at BOS Airport James
1:00-1:30 PM Luggage pickup James
1:30-2:00 PM Rent car James
2:00 PM Start turkey Sarah
2:00-3:00 PM Drive home James
2:30 PM Land at BOS Airport Emily
3:00 PM Arrive home Michael
3:00-4:00 PM Drive to airport, pick up Emily James
4:00-5:00 PM Return home with Emily James
4:30-5:00 PM Drive to Grandma’s Michael
5:00-5:30 PM Return with Grandma Michael
4:00-6:00 PM Prepare side dishes Sarah
6:00 PM Dinner served All
A.8.2 Observations on Errors
Handling Long Dependencies
Complex scheduling problems reveal cascading errors when dependencies overlap. Key constraints, especially multi-factor dependencies, often get dropped during iterative problem-solving.
Reason: Cognitive load limits simultaneous constraint tracking, making exhaustive verification difficult in single passes.
Solution Framework:
  • Isolate and enumerate atomic task dependencies.
  • Verify global constraint satisfaction.
  • Implement systematic conflict resolution.
Stale Memory and Iterative Revisions
Iterative solutions can propagate errors due to partial constraint resets.
Reason: Over-reliance on previous solutions without full constraint re-evaluation leads to compounding errors.
Relation to Gödel’s Incompleteness:
  • Systems capable of arithmetic contain unprovable truths.
  • Similarly, inherited errors hinder consistent solutions.
  • Clean-state resets necessary for error prevention.
Implementation Strategy
Reset to baseline state for each iteration, fully re-evaluating all constraints.
Core Challenges:
  • Nested dependency management.
  • Residual error prevention.
  • Cross-iteration consistency.
A.9 Experiment #2: Reactive Planner for Flight Delay
At 10:00 AM Eastern time, Sarah is notified that James’s flight is delayed by three hours, with a new arrival time of 4:00 PM. Incorporating this unexpected delay, $\mathcal{MP}$ generates a reactive plan, $\mathbf{W^{R}}$.
Early Information Agent Addition
The meta-planner adds an early information agent to monitor upstream events:
$f_{\text{early}}:\mathcal{E}_{\text{upstream}}\rightarrow\text{alerts}$ (14)
The agent’s protocol is defined as:
Table 15: Early Information Agent Specification
Component Flight Monitor Impact Analyzer
Input Flight status, departure logs, weather Alert details, workflow dependencies
Output Alert(event, severity, delay) Replan(affected_nodes, time_window)
This addition allows the workflow to initiate replanning at the earliest possible moment when upstream changes occur, significantly enhancing the system’s proactive planning capability. Since none of the planned elements have been executed, this reactive planning effectively functions as proactive planning.
In this experiment, the problem statement remains unchanged apart from James’s updated arrival time.
Initial Setup (Updated at 10:00 AM):
  • Mom (Sarah) is hosting Thanksgiving dinner at 6:00 PM in Boston. The following family members are traveling:
  • Dad (James) flying from San Francisco, landing at 4:00 PM Eastern time [UPDATED].
  • Sister (Emily) flying from Chicago, landing at 2:30 PM
  • Brother (Michael) driving from New York, estimated arrival 3:00 PM at home
  • Grandma is healthy and needs to be picked up from her home in suburban Boston
Critical Dependencies:
  • James must rent a car after landing
  • Emily must be picked up from airport, no other transportation options are allowed
  • Turkey needs 4 hours to cook, someone must be in the house once turkey is in oven for safety
  • Side dishes require 2 hours of preparation, which can overlap with turkey
  • Travel time between home and Boston airport is one hour (one-way)
  • Travel between Boston airport and grandma home is one hour (one-way)
  • Travel between home and grandma home 30 minutes (one-way)
* New Dependencies:
  • The airport luggage pickup time after landing is 30 minutes.
  • Renting a car takes 30 minutes.
  • One person can simultaneously prepare turkey and side dishes.
  • Grandma prefers Michael to pick her up, provided that it does not cause the dinner time delay.
  • Grandma and Sarah prefer not to cook together in the kitchen.
  • Traffic congestion is not factored into current planning.
Planning Question Set:
  1. 1.
    All tasks and dependencies must be strictly observed in the plan, or the plan fails.
  2. 2.
    Dinner time is strictly at 6:00 PM, all tasks must be completed by then (redundancy).
  3. 3.
    Account for the idle time of each person.
  4. 4.
    The schedule consists of three columns: time, task, and assigned person(s).
A.9.1 Results: DeepSeek Wins
None of the LLMs cannot react appropriately to this new event without clearing their context buffers. As explained in Appendix A.8.2, this limitation is evident. The key takeaway is that for future runtime frameworks, we must ensure infrastructure support for selectively invalidating stale constraints. If a workflow is already in execution, completed steps and assignments cannot be erased or altered. For example, in a stock-market investment plan, when pertinent news arrives, $\mathcal{MP}$ cannot revert completed nodes or resolved dependencies in $\mathbf{W^{R}}$. For now, we treat the reactive plan as a new plan, given that no steps have been realized in the real world by 10:00 AM.
Table 16 presents GPT4o’s plan. There are three severe constraint violations. Unfortunately, when asked to identify violations, it answers none. Therefore, $\mathcal{MP}$ is stuck without a feasible plan.
Table 16: GPT4o’s Infeasible Plan. Fail to proceed.
Time Task Assigned
10 - 2:00 PM Prep side dishes (2 hours, overlaps with turkey cooking later) Sarah
X2:30 - 3:00 PM Pick up Emily from the airport Sarah
X3:00 - 4:00 PM Prep turkey and place it in the oven (4-hour cooking time) Sarah
3:00 - 3:30 PM Michael drives to pick up Grandma Michael
3:30 - 4:00 PM Drive Grandma home Michael
4:00 - 4:30 PM James lands and gets luggage James
4:30 - 5:00 PM James rents a car James
X5:00 - 5:30 PM James drives home James
5:00 - 6:00 PM Set the table and clean kitchen Emily
5:30 - 6:00 PM Michael helps Grandma settle in and assists with final prep Michael
Table 17 depicts Claude’s plan. It violated a couple of constraints in the first two attempts, but these were minor. For instance, in the second trial, it planned for Michael’s round trip to Grandma’s home to take 30 minutes. However, the key is that Claude can recognize its own error and make corrections in the next iteration.
When asked to produce an alternate plan to reduce wait time and improve load balancing, as Michael can suffer from severe fatigue, an implicit constraint, Claude generates another feasible plan in Table 18. In this plan, James picks Emily instead of Michael. Emily has to wait for James’ availability for two hours at the airport.
There are clearly other alternatives to improve the schedule and eliminate Emily’s wait time, but none of the LLMs can figure that out. For example, using the time between 10:00 am and 1:00 pm, Sarah could have picked up Grandma and assigned her to cook, allowing Sarah to be available as a driver.
Table 17: Claude’s Reactive Plan #1, Three Iterations
Time Task Assigned
13:00 Start turkey in oven Sarah
14:00 Start side dishes Sarah
14:30 Land at airport Emily
15:00 Arrive home from NY Michael
15:00 Leave for airport Michael
16:00 Land at airport James
16:00 Pick up Emily Michael
17:00 Arrive home with Emily Michael
17:00 Leave for Grandma Michael
18:00 Arrive home with Grandma Michael
18:00 Arrive home James
18:00 Dinner starts All
Table 18: Claude’s Reactive Plan #2. Michael can rest.
Time Task Person
13:00 Start turkey Sarah
14:00 Start side dishes Sarah
14:30 Land at airport Emily
15:00 Arrive from NY Michael
16:00 Land at airport James
16:30 Leave for Grandma Michael
16:30 Get rental car James
17:00 Pick up Emily James
17:00 Pick up Grandma Michael
17:30 Return home with Grandma Michael
18:00 Arrive with Emily James
18:00 Dinner starts All
DeepSeek offers a clever alternative by routing Michael directly to the Boston airport without stopping at home first. This is a pleasant common-sense inference that the other two LLMs failed to include themselves. (This was supposed to be provided by $\mathcal{MP}$’s common-sense spatial reasoning, but it did not.)
However, Michael could drive to Grandma’s home after picking up Emily. This schedule not only saves 30 minutes but also makes Grandma happy by allowing her to surprisingly see two grandchildren at the same time.
Table 19: DeepSeek’s Reactive Plan. Three Iterations. Routining Michael directly to BOS is smart.
Time Task Assigned
10:00 AM Michael departs New York for Boston Airport (4-hour drive). Michael
2:00 PM Start cooking turkey Sarah
2:30 PM Emily lands at Boston Emily
3:00 PM Emily gets her luggage Emily
3:00 PM Michael arrives at Logan airport, picks up Emily. Michael
3:00–4:00 PM Drive Emily home Michael
4:00 PM Michael departs for Grandma Michael
4:00 PM James lands at Boston Airport James
4:00–4:30 PM James picks up luggage James
4:30–5:00 PM James rents car (30 minutes). James
4:30 PM Michael arrives at Grandma’s Michael
5:00 PM Michael & Grandma home. Grandma
5:00–6:00 PM James drives home from BOS James
4:00–6:00 PM Sarah prepares side dishes (overlaps with turkey). Sarah
6:00 PM James arrives home. Dinner served. All
A.9.2 Observations on Errors
The initial meta-planner failed to recognize a critical opportunity in early information detection. A flight delay from SFO to BOS becomes known at departure time (10:00 AM EST) rather than arrival time (1:00 PM EST). An early information agent could enable replanning three hours sooner by monitoring flight departures. To remedy this oversight, the meta-planner adds an early information agent specification, detailed in Table 15. DeepSeek was aware of this alert in a timely manner, but Claude was not.
A.10 Conclusion
Our concluding remark is that we may not be able to rely on LLMs alone to cover all constraints and react promptly to various alerts. This reinforces that the MACI architecture is on the right path to address all the aforementioned limitations of LLMs, some of which cannot be rectified.
Appendix B Validation and Recovery Protocols
The validation protocol implements a multi-stage process for ensuring state consistency. When any agent proposes a state change, the validation agent initiates a sequence of checks:
$\text{validate}(s_{t}\rightarrow s_{t+1})=\begin{cases}\text{true}&\text{if all checks pass}\\ \text{false}&\text{if any check fails}\end{cases}\vspace{-.1in}$ (15)
The protocol begins with pre-validation. Before a state transition starts, the validation agent queries relevant agents about preconditions. For a travel booking, temporal agent verifies the proposed times fit within existing schedules. Spatial agent confirms the physical feasibility of movements between locations. Role agent checks if all actors can perform their assigned functions.
During the transition, the protocol maintains atomic operations. The validation agent tracks changes across all state dimensions, ensuring partial updates cannot create inconsistent states. If the temporal agent approves a flight time but the resource agent finds insufficient seats, the entire transition fails and rolls back.
Post-validation examines the resulting state. The validation agent verifies that all constraints remain satisfied after the change. Common sense agent reviews the new state for practical issues that formal checks might miss. Strategy agent confirms the transition aligns with overall planning objectives.
When validation fails, the protocol triggers a structured recovery process:
$\text{recover}(s_{t},s_{\text{failed}})\rightarrow s_{\text{valid}}$ (16)
Recovery begins by logging the failure cause and violated constraints. The strategy agent then works with domain agents to generate alternative proposals that satisfy the constraints. This might involve relaxing non-critical constraints or exploring different approaches to meet the planning objectives.
B.1 Operations Research Techniques in Validation Protocols
The validation protocols described above align closely with established methods in operations research (OR). Some relevant techniques include:
  • Constraint Programming (CP): Focuses on solving combinatorial problems by enforcing constraints, ensuring consistency across dimensions such as temporal, spatial, and resource availability (Rossi et al., 2006).
  • Mixed-Integer Linear Programming (MILP): Optimizes decision variables subject to linear constraints and objective functions, often used in scheduling and resource allocation (Wolsey & Nemhauser, 1998).
  • Network Flow Algorithms: Validates feasibility and optimizes flows in networks by ensuring capacity, timing, and availability constraints are satisfied (Ahuja et al., 1993).
  • Dynamic Programming (DP): Decomposes problems into sequential subproblems, useful for validating processes like inter-terminal walking or luggage claiming time (Bellman, 1957).
  • Monte Carlo Simulation: Simulates scenarios to validate feasibility and robustness under uncertainty (Metropolis & Ulam, 1949).
  • Robust Optimization: Focuses on solutions that remain feasible under uncertainty, ensuring plans adapt to disruptions (Ben-Tal & Nemirovski, 2009).
Integration in Multi-Agent Systems
The validation agent also employs techniques from multi-agent systems, such as:
  • Blackboard Systems: A shared workspace for collaborative validation by different agents, ensuring global consistency (England & Engelmore, 1987).
  • Consensus Protocols: Used for distributed validation, where agents negotiate to ensure all constraints are met (Ren & Beard, 2005).
By combining these OR techniques with agent-based systems, the validation protocol ensures comprehensive and adaptive checks for workflow consistency. Future work can explore integrating heuristic methods, such as genetic algorithms or simulated annealing, to further enhance recovery processes.
Appendix C $\mathsf{MACI}$ Additional Design Considerations
C.1 Cross-Domain Generalization
While the state space dimensions—Who, Where, When, What, and Why—are broad enough to cover diverse domains, additional customization may be required for unique applications. This section examines how MACI generalizes across domains like financial planning, healthcare logistics, and supply chain optimization. The travel planning example is illustrative, emphasizing how MACI dynamically adapts state spaces and agents to domain-specific requirements.
C.2 Dynamic Agent Registration and Evolution
This section explores how agents are dynamically developed, trained, and validated for new tasks. It discusses mechanisms for evaluating new agents and integrating them into the repository without retraining the entire system, ensuring scalability and adaptability.
C.3 Scalability and Resource Efficiency
As the number of agents and task complexity grows, MACI employs strategies to manage communication overhead and optimize agent interactions. This section details techniques for clustering agents and hierarchical coordination to maintain scalability.
C.4 Empirical Evaluation Across Domains
To demonstrate MACI’s adaptability, this section presents empirical results from applying the framework to multiple domains. Examples include financial portfolio management, urban traffic planning, and hospital resource allocation, highlighting MACI’s advantages over state-of-the-art systems.
C.5 Challenges and Future Directions
While MACI addresses many limitations of LLM-based planning, challenges remain in real-time coordination, implicit knowledge integration, and robust recovery mechanisms. This section proposes future research directions to enhance MACI’s performance and applicability to novel tasks.
Appendix D Additional Tables and Figures
D.1 Notation
Table 20 presents all symbols used throughout this paper.
Table 20: Symbol Definitions
Symbol Definition Symbol Definition
Basic Sets
$\mathcal{O}$ Planning objectives $\mathcal{P}$ Available people
$\mathcal{C_{E}}$ Explicit constraints $\mathcal{C_{I}}$ Implicit constraints
$\mathcal{M}$ Performance metrics $\mathcal{Q}$ Role qualifications
Workflow Components
$\mathbf{W}$ Workflow network $\mathcal{N}$ Roles (nodes)
$\mathcal{E}$ Dependencies (edges) $\mathbf{A}$ Agent repository
$n_{i}$ Individual role $e_{ij}$ Role dependency
Agent Functions
$f_{\text{role}}$ Role-agent mapping $f_{\text{edge}}$ Edge-agent mapping
$V(\cdot)$ Validation function $\text{dist}(\cdot)$ Capability distance
$\text{map}_{\text{role}}$ Role extraction $\text{map}_{\text{edge}}$ Dependency extraction
Optimization
$A^{*}_{n}$ Selected node agents $A^{*}_{e}$ Selected edge agents
$\mathbf{W^{*}}$ Optimal workflow $V^{*}$ Best validation score
D.2 Example Common Agents
  1. 1.
    Role Manager Agent (Who): Tracks actors, their roles, and their associated constraints, ensuring that all role-based requirements are satisfied.
  2. 2.
    Spatial Agent (Where): Manages geographic and location-based constraints, verifying transitions between physical or virtual locations.
  3. 3.
    Temporal Agent (When): Handles scheduling, timing, and deadlines, ensuring alignment with temporal constraints.
  4. 4.
    Resource Agent (What): Tracks real-world resources such as tools, vehicles, or financial instruments, managing capacity, availability, and associated costs.
  5. 5.
    Reasoning and Explanation Agent (Why): Maintains the rationale behind decisions, dependencies, and alternative plans, enabling consistent alignment with objectives and providing explanations for outcomes.
  6. 6.
    Common Sense Agent: Identifies implicit constraints, integrates practical knowledge, and ensures plans align with real-world considerations.
  7. 7.
    Constraint Validation Agent: Ensures that all constraints are satisfied and that proposed plans remain feasible.
  8. 8.
    Plan Evaluation Agent: Assesses the effectiveness of plans against predefined metrics and objectives.
  9. 9.
    What-If Testing Agent: Evaluates plan robustness by simulating alternative scenarios and analyzing their impact.
  10. 10.
    Compliance and Safety Agent: Monitors adherence to safety standards, ethical principles, and regulatory frameworks.
Appendix E Traveling Salesman Problem Experiment
E.1 General Problem Specification
The TSP requires finding the shortest possible route visiting N locations exactly once, returning to the start:
  • Inputs: N locations, distance matrix D[N][N]
  • Output: Optimal tour T minimizing total distance
  • Constraints: Each location visited once, return to start
Computational Complexity - Brute Force
For N locations:
  • Number of possible tours = (N-1)!/2
  • Time complexity = O(N!)
  • Space complexity = O(N2)
Solution Methods
  1. 1.
    Exact Methods: Representative methods are Branch and Bound (Land & Doig, 1960), Dynamic Programming (Bellman, 1962), and Integer Linear Programming (Dantzig et al., 1954).
  2. 2.
    Heuristics: Methods include Nearest Neighbor, Insertion Methods, and Christofides Algorithm (3/2-approximation) (Christofides, 1976).
  3. 3.
    Meta-heuristics: This category includes Genetic Algorithms (Holland, 1992), Simulated Annealing (Kirkpatrick et al., 1983), and Ant Colony Optimization (Dorigo & Stützle, 2004).
E.2 $\mathbf{W^{*}}$: MACI Generated Planner for TSP
Node Components (N)
For TSP with n locations:
$N=\{n_{\text{route}},n_{\text{dist}},n_{\text{valid}}\},\text{ where}$ (17)
    $n_{\text{route}}$: Route generation role
    $n_{\text{dist}}$: Distance calculation role
    $n_{\text{valid}}$: Solution validation role
Edge Dependencies (E)
$E=\{e_{\text{spatial}},e_{\text{sequence}},e_{\text{complete}}\}\text{ where}$ (18)
    $e_{\text{spatial}}$: Distance constraints between locations
    $e_{\text{sequence}}$: Visit order constraints
    $e_{\text{complete}}$: Tour completion requirements
Agent Assignments
Node Agents ($A_{n}$):
  • Route Generation Agent: Generates candidate tours
  • Distance Calculator Agent: Computes tour lengths
  • Solution Validator Agent: Verifies tour validity
Edge Agents ($A_{e}$):
  • Spatial Constraint Agent: Monitors distance feasibility
  • Sequence Monitor Agent: Ensures valid visit order
  • Completion Checker Agent: Verifies tour completeness
Algorithm Selection
Based on the size of the problem, an algorithm is selected to balance performance trade-offs and mitigate the exponential computational cost of the brute-force method.
Validation Function
$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$ (19)
E.3 Experiments, From Small to Large N
  1. 1.
    N=5: Establish ground truth via brute force
  2. 2.
    N=10,20,100: Test LLM performance degradation
  3. 3.
    Metrics:
    • Solution quality vs optimal
    • Computation attempts before giving up
    • Error recognition capability
E.3.1 Small Campus Tour (N=5)
Plan an optimal route for campus tour guide visiting 5 key locations:
  • A: Admissions Office (start/end)
  • B: Library
  • C: Student Center
  • D: Science Building
  • E: Sports Complex
Distance Matrix (minutes)
$D=\begin{bmatrix}0&5&8&4&7\\ 5&0&6&3&8\\ 8&6&0&5&4\\ 4&3&5&0&6\\ 7&8&4&6&0\end{bmatrix}$ (20)
Constraints
  • Tour starts/ends at Admissions (A)
  • Each location visited exactly once
  • Total possible routes: (5-1)!/2 = 12
  • Optimal solution can be verified by hand
3.1.1. $\mathbf{W^{*}}$ Workflow Components
Node Components (N)
  • $n_{\text{route}}$: Generates permutations starting/ending at A.
  • $n_{\text{dist}}$: Computes tour length using distance matrix $D$.
  • $n_{\text{valid}}$: Checks tour validity (start/end at A, no repeats).
Edge Dependencies (E)
  • $e_{\text{spatial}}$: Enforces distance constraints from $D$.
  • $e_{\text{sequence}}$: Ensures visit order consistency.
  • $e_{\text{complete}}$: Validates all 5 locations are visited.
Agent Assignments
  • Node Agents:
    • Route Generation Agent (handles $n_{\text{route}}$)
    • Distance Calculator Agent (handles $n_{\text{dist}}$)
    • Solution Validator Agent (handles $n_{\text{valid}}$)
  • Edge Agents:
    • Spatial Constraint Agent (enforces $e_{\text{spatial}}$)
    • Sequence Monitor Agent (enforces $e_{\text{sequence}}$)
    • Completion Checker Agent (enforces $e_{\text{complete}}$)
Selected Algorithm
Brute-force.
Validation Function
$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$
3.1.2. Solution Steps
Step 1: Problem Parsing
  • Input: 5 locations with distance matrix $D$.
  • Initialize node/edge agents and constraints.
Step 2: Route Generation ($n_{\text{route}}$ Agent)
  • Generate all valid permutations: $\frac{(5-1)!}{2}=12$ routes.
  • Example permutations:
    • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
    • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
Step 3: Distance Calculation ($n_{\text{dist}}$ Agent)
  • Compute total time for each route using $D$.
Step 4: Solution Validation ($n_{\text{valid}}$ Agent)
  • Check all routes for:
    • Start/end at A (e.g., invalid route: $A\rightarrow B\rightarrow C\rightarrow D\rightarrow E\rightarrow B$).
    • No duplicate visits.
Step 5: Edge Agent Validation
  • Spatial Constraint Agent: Verify $D_{i,j}$ matches edge weights.
  • Sequence Monitor Agent: Confirm no backtracking (e.g., $B\rightarrow D$ allowed; $D\rightarrow B$ invalid unless part of loop).
Step 6: Apply Validation Function
  • Assign $V=-\infty$ to invalid routes.
[…]
  • Identify minimal $V=-24$ (i.e., maximal tour length 24 mins).
3.1.3 Solution
Optimal tour time: 242424 minutes, achieved by three routes:
  • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
  • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
  • $A\rightarrow E\rightarrow C\rightarrow B\rightarrow D\rightarrow A$
E.3.2 Large Campus Tour (N=10)
Plan an optimal route for a guided tour through 10 locations:
  • Locations: A (Admissions), B (Library), C (Student Center), …, J (Sports Complex)
  • Distance Matrix: Asymmetric travel times (minutes)
$D=\begin{pmatrix}0&12&8&15&9&14&7&11&10&6\\ 10&0&7&14&6&16&9&13&5&8\\ 9&5&0&11&8&12&10&7&15&4\\ 14&8&12&0&10&9&13&6&11&7\\ 7&13&6&9&0&8&5&12&14&10\\ 11&9&15&8&12&0&7&10&13&5\\ 5&7&10&6&11&9&0&8&12&15\\ 8&14&4&10&7&13&6&0&9&11\\ 12&6&9&7&15&10&8&5&0&14\\ 9&10&7&13&5&11&14&8&12&0\end{pmatrix}$
Table 21: Comparison of Planners and Their Performance Characteristics. $\mathcal{MP}$ provides validation to improve accuracy.
Planner Best Results Algorithm Iters. Advantages Limitations
Claude $92\rightarrow 66$ mins Nearest Neighbor 3 Efficiently implements greedy heuristic approach Makes data reading errors, compromising solution accuracy
GPT4o $75\rightarrow 68$ mins Genetic 3 Identifies effective termination conditions Unable to implement exact algorithms like Held-Karp
DeepSeek $60$ mins Held-Karp 1 Implements optimal algorithm correctly None observed for this problem size
$\mathcal{MP}$ + Claude $66\rightarrow 60$ mins Ant Colony Optimization 2 Provides validation and suggests iteration increases for improvement Requires external guidance for algorithm selection and parameter tuning
MP + GPT4o $60$ mins Ant Colony Optimization 1 Achieves optimal solution with precise execution Requires more computational resources with larger ant population and iteration count
MP + DeepSeek $60$ mins Ant Colony Optimization 1 Combines efficient algorithm selection with optimal parameter tuning None significant for given problem
Algorithm Selection
Based on the size of the problem, $\mathcal{MP}$ selected the Ant Colony Optimization (Dorigo & Stützle, 2004) algorithm to achieve at least a 4x speedup. For $N\geq 10$, an approximate method is recommended.
3.2.1. ACO Method
Parameters
  • 100 ants, 50 iterations, and $\rho=0.1$ evaporation
  • $\alpha=1$ (pheromone weight), and $\beta=2$ (heuristic weight)
The termination criteria can be modified to stop the algorithm if no meaningful improvement is observed after $k$ consecutive iterations.
Algorithm
1:  Initialize $\tau_{ij}\leftarrow 1.0$, $\eta_{ij}\leftarrow 1/D_{ij}$
2:  for 50 iterations do
3:     for all 100 ants do
4:        Build tour using $P_{ij}=\frac{[\tau_{ij}]^{1}[\eta_{ij}]^{2}}{\sum[\tau_{ik}]^{1}[\eta_{ik}]^{2}}$
5:        Record tour length $L_{k}$
6:     end for
7:     Evaporate pheromones: $\tau_{ij}\leftarrow 0.9\tau_{ij}$
8:     Deposit pheromones: $\tau_{ij}\leftarrow\tau_{ij}+\sum\frac{10}{L_{k}}$
9:     Track best tour
10:  end for
3.2.2. Performance Comparison
The goalpost is the optimal time of 606060 minutes. Table 21 compares six different configurations, and three of the six achieve the optimal answer. Although TSP is a relatively simple scheduling problem with just one actor and no parallel execution, the benefit of having $\mathcal{MP}$ to validate results is still helpful to Glaude and GPT4o.
When asked to solve the problem without $\mathcal{MP}$, Glaude and GPT4o initially chose brute force, then switched to an approximation method without thorough deliberation (or perhaps they did, but did not output their reasoning process). However, DeekSeek picked Held-Karp, a computationally expensive method, even more expensive than brute force, arguing that the absolute computation time for $N=10$ is only $0.2$ seconds. $\mathcal{MP}$ was more deliberate, opting for brute force when $N=5$ and ACO for $N=10$.
E.4 TSP Experiment Conclusion
This simple task demonstrates that $\mathcal{MP}$ can be valuable for monitoring the execution process, validating the correctness of intermediate results, and suggesting more efficient algorithmic approaches.